a chemist wants to produce a 30% alcohol solution. how many liters of 20% and 50% solution should be mixed in order to get 12 liters of 30% solution?
Answer by mananth(16946)

--------- percent ----------------quantity
Orange Juice I 20 ----------------x L
Orange Juice II 50 ----------------12-x L
Total 30.00% ---------------- 12

20x+50(12-x)=30 *12
20x+600 -50x=360
20x-50x =360-600
-30x=-240
/-30
x=8 L 20.00%
4 L 50.00%